# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 6.2. The action of $s$ mimics a $\kappa$-length ride on the $S_{\rho}$-flow
Recall the specification of $S_{\rho}(x,\cdot)$ as an ODE solution from Definition 1.2.
Lemma 6.2 determines the linear coefficient in the small-$\kappa$ expansion of $\phi_{0}$ and $\phi_{1}$ and controls the $O(\kappa^{2})$ term. It thereby offers a counterpart estimate for the map $s:\phi_{0}\mapsto\phi_{1}$. To stand ready for a comparison of the $s$-iterates with the flow $S_{\rho}$, it also presents a time-$\kappa$ evolution estimate for $S_{\rho}$.
@@
Let $(\rho,\beta)\in(0,\infty)^{2}$ and suppose that $\kappa\in\big(0,(1+\rho)^{-1}/2\big)$.
@@
    Further suppose that $\rho^{2}\kappa\leq 1$. Then
    $s(x)\,=\,x-\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\,+\,x\,\Theta_{3}(\kappa,\rho,\beta)\kappa^{2}$
    with $|\Theta_{3}(\kappa,\rho,\beta)|\leq 52\rho(1+\rho)^{3}$.
@@
    For $x\in(0,\infty)$ and $\kappa>0$,
    $S_{\rho}(x,\kappa)\,=\,x-\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\,+\,x\,\Theta_{4}(\kappa,\rho,x)\,\kappa^{2}$
    with $|\Theta_{4}(\kappa,\rho,x)|\leq 64\rho^{2}(1+\rho)$.
Proof: (1). We may express $\phi_{0}=\beta N/D$, with
$N=(1-\kappa)\beta^{2\rho}+2(1+\rho\kappa)\beta^{\rho}+1+\kappa,\quad D=(1-\kappa)\beta^{2\rho}+2(1-\rho\kappa)\beta^{\rho}+1+\kappa.$
@@
$R(\kappa,\beta)\,\leq\,2\frac{|D_{1}|\cdot|D_{1}-N_{1}|}{D_{0}^{2}}\,\leq\,2(1+\rho)$
@@
(2). By Lemma 2.3(1,2), $\phi_{0}$ and $\phi_{1}$ are increasing bijections of $(0,\infty)$ under our hypothesis; thus, the map $s:(0,\infty)\to(0,\infty)$, which by definition sends $\phi_{0}$ to $\phi_{1}$, is well defined.
@@
$s(x)=x\,-\,\frac{8\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}\,\kappa\;+\;\beta\,\Theta(\kappa,\rho,\beta)\kappa^{2}\,$,
But $\frac{8\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}=\frac{8\rho x^{1+\rho}}{(1+x^{\rho})^{2}}+R$ where $|R|\leq 8\rho|\beta-x|D$, with $D=\sup\big\{\big|\tfrac{{\rm d}}{{\rm d}z}\tfrac{z^{1+\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. By Lemma 6.2(1),
@@
$s(x)$ $=$ $x\,-\,\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;8\rho D\bigg(\frac{4\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}\,\kappa\;+\;\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\bigg)\kappa\,+\,\beta\,\Theta(\kappa,\rho,\beta)\kappa^{2}$
$=$ $x-\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\,+\,\beta\,\Theta_{3}(\kappa,\rho,\beta)\kappa^{2}$
where (since $\kappa\leq 1$)
$|\Theta_{3}|\leq|\Theta|+32\rho^{2}DD_{0}+8\rho D|\Theta_{1}|$
with $D_{0}=\sup\big\{\big|\tfrac{z^{\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. Since $D_{0}\leq 1/4$ and $D\leq 2(1+\rho)$, the latter right-hand side is at most
$4\rho(1+\rho)+32\rho^{2}DD_{0}+16D\rho^{2}(1+\rho)\leq 4\rho(1+\rho)+16\rho^{2}(1+\rho)+32\rho^{2}(1+\rho)^{2}\leq 52\rho(1+\rho)^{3}\,.$
The expression $x=\phi_{0}(\kappa,\rho,\beta)$ as specified in (10) is at least $\beta$: indeed, the right-hand numerator there is clearly positive, and so is the denominator under the hypothesis that $\rho^{2}\kappa\leq 1$, as we noted in the proof of Lemma 2.3(3); the bracketed term in the numerator is clearly less than the numerator, so the expression overall is at least $\beta$. This bound permits us to replace the prefactor of $\beta$ by $x$ in the last right-hand term in the above expression for $s(x)$. Thus we obtain Lemma 6.2(2).
(3). Writing $f(x)=-\frac{8\rho x^{1+\rho}}{(1+x^{\rho})^{2}}$, recall that $S_{\rho}(x,u)$ solves $\frac{{\rm d}}{{\rm d}u}S_{\rho}(x,u)=f(S_{\rho}(x,u))$ with $S_{\rho}(x,0)=x$. Observe that $f(x)<0$ for all $x>0$, so $S_{\rho}(x,u)$ is decreasing in $u$ and satisfies $0<S_{\rho}(x,u)\leq x$ for all $u\geq 0$. Since $f$ is Lipschitz, the Picard-Lindelöf theorem [10, Theorem $I.3.1$] implies that $S_{\rho}(x,u)$ has the integral form
$S_{\rho}(x,\kappa)=x+\int_{0}^{\kappa}f(S_{\rho}(x,u))\,du\,.$
Since $S_{\rho}(x,u)\leq x$, we have for $u\in[0,\kappa]$,
$|S_{\rho}(x,u)-x|=\left|\int_{0}^{u}f(S_{\rho}(x,v))\,{\rm d}v\right|\leq u\,\sup_{z\in[0,x]}|f(z)|\leq\kappa\cdot 8\rho x\,.$
By the mean-value theorem, $f(S_{\rho}(x,u))=f(x)+f^{\prime}(\xi_{u})\,(S_{\rho}(x,u)-x)$ for some $\xi_{u}\in[S_{\rho}(x,u),x]\subset[0,x]$. Integrating, the remainder is
$R(x,\kappa):=S_{\rho}(x,\kappa)-x-\kappa f(x)=\int_{0}^{\kappa}f^{\prime}(\xi_{u})\big(S_{\rho}(x,u)-x\big)\,{\rm d}u\,.$
Differentiating $f$, we readily see that $|f^{\prime}(z)|\leq 8\rho(1+\rho)/(1+x^{\rho})^{2}$. Using the preceding bound on $|S_{\rho}(x,u)-x|$ alongside $|f^{\prime}(z)|\leq 8\rho(1+\rho)$, we find that
$|R(x,\kappa)|\leq\int_{0}^{\kappa}|f^{\prime}(\xi_{u})|\,|S_{\rho}(x,u)-x|\,{\rm d}u\leq\kappa\cdot 8\rho(1+\rho)\cdot(\kappa\cdot 8\rho x)=64\rho^{2}(1+\rho)x\,\kappa^{2}\,.$
Then setting $\Theta_{4}(\kappa,\rho,x)=R_{2}(x,\kappa)/(x\kappa^{2})$, we obtain
$S_{\rho}(x,\kappa)=x+\kappa f(x)+x\,\Theta_{4}(\kappa,\rho,x)\,\kappa^{2}\,$,
with $\Theta_{4}(\kappa,\rho,x)|\leq 64\rho^{2}(1+\rho)$, which completes the proof. ∎
@@ 6.3. The scaled $s$-orbit tracks the $S_{\rho}$-flow
Let $\rho\in(0,1]$ and $x\in D$. As functions of the argument $\bullet\in\mathbb{R}$, we have that
$s_{\lfloor\kappa^{-1}\bullet\rfloor}(x)\,\stackrel{{\scriptstyle\kappa}}{{\simeq}}\,S_{\rho}(x,\bullet)\,.$
Remark. We may also take $\rho\geq 1$ provided that $(\kappa,\rho)$ lies in $W$ as specified in (8) and the $\stackrel{{\scriptstyle\kappa}}{{\simeq}}$ notation is modified to permit $\rho$-dependent constants.
@@
Developing a concept from Section 2.1, we say that a bijection $h:(0,\infty)\to(0,\infty)$ is role-reversal symmetric if its inverse satisfies $h^{-1}(x)=1/h(1/x)$. Indeed, Proposition 2.1 shows this property for $s$. The flow satisfies it also: by Proposition 5.4(1), $x\to S_{\rho}(x,r)$ is role-reversal symmetric for any $r>0$.
By Lemma 6.2(2,3) (and $\rho\leq 1$), the maps from $(0,\infty)$ to $(0,\infty)$ given by $x\to s(x)$ and $x\to S_{\rho}(x,\kappa)$ belong to $\mathcal{I}_{\rho}(\kappa,C)$ with $C=500$.
@@
    $\big|h_{i+1}(x)-h_{i+1}^{\prime}(x)\big|\leq(1+8C^{\prime}\rho\kappa)\big|h_{i}(x)-h_{i}^{\prime}(x)\big|+1000\,\kappa^{2}\max\big\{|h_{i}(x)|,|h^{\prime}_{i}(x)|\big\}\,.$
    where $C^{\prime}$ denotes the supremum (which is readily seen to be finite) of the absolute value of the derivative of $\frac{z^{1+\rho}}{(1+z^{\rho})^{2}}$ over $(0,\infty)$.
@@
To confirm that the proposition follows from the lemma, take $h(\bullet)=s(\bullet)$ and $h^{\prime}=S_{\rho}(x,\bullet)$. For $r\in\mathbb{R}$, set $i=\lfloor\kappa^{-1}r\rfloor$. Since $x$ lies in the bounded central domain $D$ and $\rho\leq 1$, Lemma 6.4(2) implies that
$\lvert h_{i}(x)-h_{i}^{\prime}(x)\rvert\,\leq\,2C\kappa\,e^{C_{2}r}\,\max\big\{|h_{i}(x)|,|h^{\prime}_{i}(x)|\big\}$
@@
Since $\psi\leq 0$ establishes IH($i+1$), the inductive argument will be complete provided that we show the above right-hand side is at most zero, for which it suffices to prove a bound of the form $\tfrac{\omega_{i+1}(x)}{\omega_{i}(x)}\geq 1-B\kappa\rho$ alongside
$C\big(1+8C^{\prime}\rho\kappa-(1+C_{2}C^{\prime}\kappa x)(1-B\kappa)\big)+C_{1}\kappa\,e^{-C_{2}C^{\prime}\kappa xi}\leq 0\,.$ (60)
We justify the lower bound on the ratio $\omega_{i+1}(x)/\omega_{i}(x)$ as follows. By Lemma 6.2(2), for each iterate we have
$h_{i+1}(x)=h_{i}(x)-\frac{8\rho\,h_{i}(x)^{1+\rho}}{(1+h_{i}(x)^{\rho})^{2}}\,\kappa+h_{i}(x)\,\Theta_{3}(\kappa,\rho,\beta)\kappa^{2},$
and similarly for $h^{\prime}_{i}(x)$. Here note that since $|\Theta_{3}|\leq 52\rho(1+\rho)^{3}\leq 500\rho$, we have $\kappa|\Theta_{3}|\leq\rho$ in view of $\kappa$ being less than a small positive constant. Since $z^{\rho}/(1+z^{\rho})^{2}$ is bounded, $|h_{i+1}(x)|\;\geq\;|h_{i}(x)|\big(1-B\kappa\rho\big)$ and likewise for the $h^{\prime}$-sequence; taking the maximum, we obtain the claimed lower bound on $\omega_{i+1}(x)/\omega_{i}(x)$.
To obtain (60), note that its left-hand side is at most

### v4

@@ 6.2. The action of $s$ mimics a $\kappa$-length ride on the $S_{\rho}$-flow
In the discrete context, the map $s:\phi_{0}\mapsto\phi_{1}$ sends the central ratio of any ABMN($\kappa,\rho$) element to the counterpart quantity associated with the left shift of this element by one unit. The integer lattice is scaled by a factor of $\kappa$ in the transition from the discrete to the continuum. Counterpart to the $s$-argument $x=\phi_{0}\in(0,\infty)$ in the scaled setting is the flow index $\xi\in(0,\infty)$, as noted after Definition 1.20. Recalling the flow $S_{\rho}(\xi,\bullet)$, specified as the solution of an ODE in Definition 1.14, we regard the time-$\kappa$ evolution $\xi=S_{\rho}(\xi,0)\mapsto S_{\rho}(\xi,\kappa)$ as counterpart to a single application of the map $s$.
The purpose of this section is to make this correspondence quantitative. Lemma 6.2 provides a sharp single-step comparison between the discrete map $s$ and the time-$\kappa$ evolution of the flow $S_{\rho}$, showing that these two operations are close to each other.
More precisely, the lemma determines the linear coefficient in the small-$\kappa$ expansions of $\phi_{0}$ and $\phi_{1}$ and controls the $O(\kappa^{2})$ terms. It thereby yields such an estimate for the map $s:\phi_{0}\mapsto\phi_{1}$; and it furnishes a counterpart for the time-$\kappa$ evolution of $S_{\rho}$. This prepares for a comparison of iterates of $s$ with progression along the $S_{\rho}$-flow.
The choice of domain variables $x$ and $\xi$ in Lemma 6.2(2,3) is characteristic of the discrete and continuous contexts. Setting these variables equal serves to make explicit the approximation discussed above.
@@
Let $(\rho,\beta)\in(0,\infty)^{2}$ and suppose that $\kappa>0$ is less than $\min\big\{\rho^{-2},(1+\rho)^{-1}/2\big\}$.
@@
    We also have that
    $s(x)\,=\,x-\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\,+\,x\,\Theta_{3}(\kappa,\rho,x)\kappa^{2}$
    with $|\Theta_{3}(\kappa,\rho,x)|\leq 52\rho(1+\rho)^{3}$.
@@
    For $\xi\in(0,\infty)$ and $\kappa>0$,
    $S_{\rho}(\xi,\kappa)\,=\,\xi-\frac{8\rho\,\xi^{1+\rho}}{(1+\xi^{\rho})^{2}}\,\kappa\,+\,\xi\,\Theta_{4}(\kappa,\rho,\xi)\,\kappa^{2}$
    with $|\Theta_{4}(\kappa,\rho,\xi)|\leq 64\rho^{2}(1+\rho)$.
Proof. Recall that, for the given value of the pair $(\kappa,\rho)$, $\phi_{0}$ and $\phi_{1}$ are functions of $\beta\in(0,\infty)$ specified by the formulas in Definition 1.7. The variable $x=\phi_{0}$ is the input for the map $s$. Our hypotheses ensure that the $(\kappa,\rho)$-pair lies in the region $W$ specified by $\rho^{2}\kappa\leq 1$ in (1.7), so that Lemma 2.1(2) implies that the relationship $x\longleftrightarrow\beta$ is bijective on $(0,\infty)$.
(1). We may express $\phi_{0}=\beta N/D$, with
$N=(1-\kappa)\beta^{2\rho}+2(1+\rho\kappa)\beta^{\rho}+1+\kappa,\quad D=(1-\kappa)\beta^{2\rho}+2(1-\rho\kappa)\beta^{\rho}+1+\kappa.$ (6.8)
@@
$R(\kappa,\beta)\,\leq\,2\frac{|D_{1}|\cdot|D_{1}-N_{1}|}{D_{0}^{2}}\,\leq\,2\rho(1+\rho)$
@@
(2). As we noted at the outset of the proof, $\phi_{0}$ and $\phi_{1}$ are increasing bijections of $(0,\infty)$ under our hypothesis; thus, the map $s:(0,\infty)\to(0,\infty)$, which by definition sends $\phi_{0}$ to $\phi_{1}$, is well defined.
@@
$s(x)=x\,-\,\frac{8\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}\,\kappa\;+\;\beta\,\Theta(\kappa,\rho,\beta)\kappa^{2}\,.$
But $\frac{8\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}=\frac{8\rho x^{1+\rho}}{(1+x^{\rho})^{2}}+R$ where $|R|\leq 8\rho|\beta-x|Q$, with $Q=\sup\big\{\big|\tfrac{{\rm d}}{{\rm d}z}\tfrac{z^{1+\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. By Lemma 6.2(1),
@@
$s(x)$ $=$ $x\,-\,\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\;+\;8\rho Q\bigg(\frac{4\rho\,\beta^{1+\rho}}{(1+\beta^{\rho})^{2}}\,\kappa\;+\;\beta\,|\Theta_{1}(\kappa,\rho,\beta)|\kappa^{2}\bigg)\kappa\,+\,\beta\,\Theta(\kappa,\rho,\beta)\kappa^{2}$
$=$ $x-\frac{8\rho\,x^{1+\rho}}{(1+x^{\rho})^{2}}\,\kappa\,+\,\beta\,\Theta_{3}(\kappa,\rho,x)\kappa^{2}$
where we are able to take $\Theta_{3}$ to be a function of $x$, rather than of $\beta$, because of the bijective relationship between the two variables; since $\kappa\leq 1$, we have
$|\Theta_{3}|\leq|\Theta|+32\rho^{2}QQ_{0}+8\rho Q|\Theta_{1}|$
with $Q_{0}=\sup\big\{\big|\tfrac{z^{\rho}}{(1+z^{\rho})^{2}}\big|:z\in(0,\infty)\big\}$. Since $Q_{0}\leq 1/4$ and $Q\leq 2(1+\rho)$, the latter right-hand side is at most
$4\rho(1+\rho)+32\rho^{2}QQ_{0}+16Q\rho^{2}(1+\rho)\leq 4\rho(1+\rho)+16\rho^{2}(1+\rho)+32\rho^{2}(1+\rho)^{2}\leq 52\rho(1+\rho)^{3}\,.$
In the notation from (6.8) $x=\phi_{0}=\beta N/D$ with $D>0$ and $N-D=4\rho\kappa\beta^{\rho}\geq 0$. Thus, $x\geq\beta$. This bound permits us to replace the prefactor of $\beta$ by $x$ in the last right-hand term in the above expression for $s(x)$. Thus we obtain Lemma 6.2(2).
(3). Writing $f(z)=-\frac{8\rho z^{1+\rho}}{(1+z^{\rho})^{2}}$, recall that $S_{\rho}(\xi,u)$ solves $\frac{{\rm d}}{{\rm d}u}S_{\rho}(\xi,u)=f(S_{\rho}(\xi,u))$ with $S_{\rho}(\xi,0)=\xi$. Observe that $f(z)<0$ for all $z>0$, so $S_{\rho}(\xi,u)$ is decreasing in $u$ and satisfies $0<S_{\rho}(\xi,u)\leq\xi$ for all $u\geq 0$. Since $f$ is Lipschitz, the Picard-Lindelöf theorem [9, Theorem $I.3.1$] implies that $S_{\rho}(\xi,u)$ has the integral form
$S_{\rho}(\xi,\kappa)=\xi+\int_{0}^{\kappa}f(S_{\rho}(\xi,r))\,dr\,.$
Since $S_{\rho}(\xi,u)\leq\xi$, we have for $u\in[0,\kappa]$,
$|S_{\rho}(\xi,u)-\xi|=\left|\int_{0}^{u}f(S_{\rho}(\xi,v))\,{\rm d}v\right|\leq u\,\sup_{z\in[0,\xi]}|f(z)|\leq\kappa\cdot 8\rho\xi\,.$
By the mean-value theorem, $f(S_{\rho}(\xi,u))=f(\xi)+f^{\prime}(\varphi_{u})\,(S_{\rho}(\xi,u)-\xi)$ for some $\varphi_{u}\in[S_{\rho}(\xi,u),\xi]\subset[0,\xi]$. Integrating, the remainder is
$R(\xi,\kappa):=S_{\rho}(\xi,\kappa)-\xi-\kappa f(\xi)=\int_{0}^{\kappa}f^{\prime}(\varphi_{u})\big(S_{\rho}(\xi,r)-\xi\big)\,{\rm d}r\,.$
Differentiating $f$, we readily see that $|f^{\prime}(z)|\leq 8\rho(1+\rho)/(1+z^{\rho})^{2}$. Using the preceding bound on $|S_{\rho}(\xi,u)-\xi|$ alongside $|f^{\prime}(z)|\leq 8\rho(1+\rho)$, we find that
$|R(\xi,\kappa)|\leq\int_{0}^{\kappa}|f^{\prime}(\varphi_{r})|\,|S_{\rho}(\xi,r)-\xi|\,{\rm d}r\leq\kappa\cdot 8\rho(1+\rho)\cdot(\kappa\cdot 8\rho\xi)=64\rho^{2}(1+\rho)\xi\,\kappa^{2}\,.$
Then setting $\Theta_{4}(\kappa,\rho,\xi)=R(\xi,\kappa)/(\xi\kappa^{2})$, we obtain
$S_{\rho}(\xi,\kappa)=\xi+\kappa f(\xi)+\xi\,\Theta_{4}(\kappa,\rho,\xi)\,\kappa^{2}\,$,
with $\Theta_{4}(\kappa,\rho,\xi)|\leq 64\rho^{2}(1+\rho)$, which completes the proof. ∎
@@ 6.3. The scaled $s$-orbit tracks the $S_{\rho}$-flow
Let $\rho\in(0,1]$ and $\xi\in D$. As functions of the argument $\bullet\in\mathbb{R}$, we have that
$s_{\lfloor\kappa^{-1}\bullet\rfloor}(\xi)\,\stackrel{{\scriptstyle\kappa}}{{\simeq}}\,S_{\rho}(\xi,\bullet)\,.$
Our usual notation for the variable in the map $s$ and its iterates is $x$. It is taken to be $\xi$ here because the scaled $s$-iterate is being directly compared to the flow $S_{\rho}(\xi,\bullet)$ whose index $\xi$ governs the choice of notation. We also mention that $\rho\geq 1$ may be taken in the above result, provided that $(\kappa,\rho)$ lies in $W$ as specified in (1.7) and the $\stackrel{{\scriptstyle\kappa}}{{\simeq}}$ notation is modified to permit $\rho$-dependent constants.
@@
Developing a concept from Section 2.2, we say that a bijection $h:(0,\infty)\to(0,\infty)$ is role-reversal symmetric if its inverse satisfies $h^{-1}(x)=1/h(1/x)$. Indeed, Proposition 2.5 shows this property for $s$. The flow satisfies it also: by Proposition 5.4(1), $\xi\to S_{\rho}(\xi,u)$ is role-reversal symmetric for any $u>0$.
By Lemma 6.2(2,3) (and $\rho\leq 1$), the maps from $(0,\infty)$ to $(0,\infty)$ given by $x\to s(x)$ and $\xi\to S_{\rho}(\xi,\kappa)$ belong to $\mathcal{I}_{\rho}(\kappa,C)$ with $C=500$.
@@
    $\big|h_{i+1}(x)-h_{i+1}^{\prime}(x)\big|\leq(1+8C^{\prime}\rho\kappa)\big|h_{i}(x)-h_{i}^{\prime}(x)\big|+1000\,\kappa^{2}\max\big\{|h_{i}(x)|,|h^{\prime}_{i}(x)|\big\}$
    for $x\in(0,\infty)$. Here $C^{\prime}$ denotes the supremum (which is readily seen to be finite) of the absolute value of the derivative of $\frac{z^{1+\rho}}{(1+z^{\rho})^{2}}$ over $(0,\infty)$.
@@
To confirm that the proposition follows from the lemma, take $h(\bullet)=s(\bullet)$ and $h^{\prime}(\bullet)=S_{\rho}(\bullet,\kappa)$. For $u\in\mathbb{R}$, set $i=\lfloor\kappa^{-1}u\rfloor$. Since $\xi$ lies in the bounded central domain $D$ and $\rho\leq 1$, Lemma 6.4(2) implies that
$\lvert h_{i}(\xi)-h_{i}^{\prime}(\xi)\rvert\,\leq\,2C\kappa\,e^{C_{2}u}\,\max\big\{|h_{i}(\xi)|,|h^{\prime}_{i}(\xi)|\big\}$
@@
Since $\psi\leq 0$ establishes IH($i+1$), the inductive argument will be complete provided that we show the above right-hand side is at most zero, for which it suffices to prove a bound of the form $\tfrac{\omega_{i+1}(x)}{\omega_{i}(x)}\geq 1-B\kappa\rho$ with $B=B(C)$ alongside
$C\big(1+8C^{\prime}\rho\kappa-(1+C_{2}C^{\prime}\kappa x)(1-B\kappa)\big)+C_{1}\kappa\,e^{-C_{2}C^{\prime}\kappa xi}\leq 0\,.$ (6.9)
We justify the lower bound on the ratio $\omega_{i+1}(x)/\omega_{i}(x)$ as follows. Since $h\in\mathcal{I}_{\rho}(\kappa,C)$, each iterate satisfies
$h_{i+1}(x)=h_{i}(x)-\frac{8\rho\,h_{i}(x)^{1+\rho}}{(1+h_{i}(x)^{\rho})^{2}}\,\kappa+h_{i}(x)\,O(1)\,\kappa^{2}$
with $|O(1)|\leq C$; and similarly for $h^{\prime}_{i}(x)$. Since $\frac{z^{\rho}}{(1+z^{\rho})^{2}}$ is bounded on $(0,\infty)$ and $\kappa$ is taken sufficiently small, it follows that
$|h_{i+1}(x)|\geq|h_{i}(x)|\big(1-B\rho\kappa\big),$
and likewise for the $h^{\prime}$-sequence. Taking the maximum yields the claimed lower bound on $\omega_{i+1}(x)/\omega_{i}(x)$.
To obtain (6.9), note that its left-hand side is at most